A particle on a stretched string supporting a travelling wave, takes 5.0 ms to move from Its mean position to the extreme position. The distance between two consecutive particles, which are at their
mean positions, is 2.0 cm. Find the frequency, the wavelength and the wave speed.

Grade:12

2 Answers

Arun
6 years ago
Tr
Arun
6 years ago
Time period, T = 4 * 5 ms = 20 * 10–3 = 2 * 10–2 s λ = 2 * 2 cm = 4 cm frequency, f = 1/T = 1/(2 * 10–2) = 50 s–1 = 50 Hz Wave speed = λ f = 4 * 50 m/s = 2000 m/s = 2 m/s